OpenStudy (anonymous):

How do I solve this integral: 1/(zsqrt(4+ln(z)))dz

OpenStudy (irishboy123):

\[\int\limits \frac{1}{z \sqrt{4+lnz}}dz = \int\limits \frac{1}{z} (4+lnz)^{-\frac{1}{2}}dz\] second way shows that it can be done by inspection, or you can use a sub let u = 4+lnz, du = dz/z, and do it that way.